Two opposite angles of a parallelogram have measures 3x-20 and 2(x+5) respectively. Find the measures of the angles
Two consecutive angles of a parallelogram have measures 4(x+5) and 2(3x+20) respectively. Find the measures of the angles.
Two consecutive angles of a parallelogram have measures 2(x+12) and 2(5x+6) respectively. Find the measures of the angles.

Answer by stanbon(75887)

Two opposite angles of a parallelogram have measures 3x-20 and 2(x+5) respectively. Find the measures of the angles.
---
Opposite angles are equal.
3x-20 = 2(x+5)
3x-20 = 2x+10
x = 30
---
1st angle = 3x-20 = 3*30-20 = 70
2nd angle = 2x+10 = 2*30+10 = 70
===========================================
Two consecutive angles of a parallelogram have measures 4(x+5) and 2(3x+20) respectively. Find the measures of the angles.
Consecutive angles are supplementary.
4(x+5) + 2(3x+20) = 180
2(x+5) + (3x+20) = 90
5x + 30 = 90
x = 12
---
1st angle = 4(12+5) = 68
2nd angle = 2(3*12+20) = 112
=============================================
Cheers,
Stan H.